### v1

@@ 2.5 Inauthentic Activity
With any social media dataset, understanding human behaviour through it also requires understanding how much of the data is produced through inauthentic activity [28]. Given that there are no publicly available bot detection tools for TikTok, we used simple heuristics for bot detection [29]. We find that there are 46 accounts with more than 100 videos each posted in the hour we looked at, and the top 5 highest frequency posting accounts, with between 1500 and 6000 videos in the hour we looked at, all have the default username of user followed by 9 random digits. Their videos have an average of 1.9 views each, but account for 0.28% of the total 1 hour dataset. Looking at comments, there are 831 accounts that have at least 10 comments and whose mean commenting frequency is higher than 1 every 3 seconds, with most of their comments (88%) consisting only of emojis. It’s clear from these initial statistics that there is automated activity on the platform, but these methods can only produce a rough lower bound on the extent of automated activity on the platform. We therefore hope future work can narrow our bounds on the amount of inauthentic activity on the platform and in this dataset.
@@ 3 Discussion
The data we report in this work is critical for calibrating our understanding of work on TikTok. For example, when we look at the invasion of Ukraine on TikTok [30], understanding that no TikTok posts are tagged with a Russian origin, and that posts from Ukraine have an average view count (4.1k) of approximately half way between the global average (2.2k) and that of the USA (6.3k) is useful context. When looking at the presence of the global south on TikTok [22], learning that posts from countries in the global south [31] account for 81% of the video posts, while only obtaining an average view count of 1.5k adds scale to our understanding.
@@
For work on Palestinian activism on TikTok [32], we can show that 0.02% of the videos in our hour dataset were classified as a Palestine protest associated topic, but those videos had a significantly lower (p < 0.02) average view count of 440 views compared to the hour dataset view average of 2.5k. For work looking at the role of war on TikTok[33], we can show that 0.15% of the videos in our hour dataset were classified as a military/soldier associated topic, and those videos received an average of 5.2k views. And for work looking at the link between TikTok and public health [34], we can add context by showing that videos associated with the topic ‘Medical Examination and Healthcare Setting’ had a significantly higher (p < 0.02) average view mean of 28.8k views than the average.
@@
In order to allow others to use the data we have collected here to calibrate their studies, we have released aggregate data distributions. We understand that we won’t cover every use case with this data release, so we invite interested parties to contact us for additional specific statistics.
@@
Our dataset is not perfect. Collecting the data for this dataset took 5 months, meaning that the first batch of data was collected 2 months after the day chosen, and the last batch of data was collected 7 months after the day chosen. This introduces temporal bias into our dataset, where older videos will accumulate more engagement, and are also more likely to be taken down, or deleted. However, looking at our comment engagement offset (Fig. 6), we can see that the median time of engagement for all comment authors is around 29.9 hours after the video was posted. It seems likely that the median time from video post to comment engagement is similar to the median time offset for other engagement, meaning that at least for engagement, the temporal bias should have minimal effect. We can see from Figure 2(b) that there is a decrease in video count over the hour we collected down to 83.7% that of the beginning of the hour, which could be due to temporal collection bias, true post volume decrease, or other factors.
@@
We also cannot speak to the proportion of authentic to inauthentic content in this dataset. Detecting bots and influence operations is a pressing issue on TikTok, but is still difficult. We know that bots exist on the platform [35], but it is unlikely that existing bot detection methods transfer well [13], due to the difference in text distribution and platform properties. In this dataset, we do not have full chronological data timelines to allow methods that exploit timestamps to work fully [36], or network data to allow graph methods to work [37]. While we used simple heuristics to examine the least sophisticated types of bots on the platform, in the age of LLMs and advanced influence operations, more comprehensive methods must be used to be able to detect this activity, which we will leave for future work.
Having said this, we believe the statistics we have presented here to be of high value to the community. With this work we have presented the first large vertical slice of posts from TikTok, and delivered the essential statistics to help understand the platform. Our data release will allow those doing studies of TikTok to contextualize their findings, and work to build on understanding of this global platform.
@@ 4.1 ID Generation
We have a starting point, which is that we know that if we convert the integer to a 64 bit binary representation, we can take the first 32 bits, convert these 32 bits back to an integer representation, we get a unix timestamp, which tends to be within seconds of the recorded creation time of the video recorded in the video metadata [38]. This indicates that within the total 64 bits of the ID, there is encoded information. Many large scale distributed systems that need to generate globally unique IDs, use systems similar to the Snowflake system [39, 40, 41, 42]. If the ID generation method is similar to snowflake, then we would expect that the rest of the bits are some combination of milliseconds data, and data relating to TikTok’s infrastructure (worker IDs, sharding IDs, sequence IDs etc.). If we can understand what composes the rest of these bits, when generating IDs, we can reduce the search space from $4.3\times 10^{9}$ combinations down to a manageable number.
@@
Canada Seedlist 118994 99.93%
Germany Seedlist 35256 99.94%
Nigeria Country Name Search 41127 99.86%
Indonesia Country Name Search 19509 99.83%
Brazil Country Name Search 14079 99.77%
All 225127 99.97%
@@
The Good-Turing estimator gives us confidence in what percentage of videos we have for each geographical area, but the question remains if we have good geographic coverage. After looking at the geographic distribution of the videos collected, we saw that the notable places where video counts were lower than others were East Asia, Oceania, India, and Russia. Though there are good reasons to believe that video counts in India and Russia would naturally be lower due to the ban of TikTok in those countries [23, 48], we collected extra datasets correlated with those regions (India, Australia, Japan, Russia) to check if there were previously undiscovered IDs, see details in Table 4. We found that the unique IDs we had previously collected covered at least 90% of the videos from those areas, and for India and Japan, at least 95%, which we feel is sufficient to ensure a dataset is representative of those areas.
@@
Russia Россия 10288 99.35% 91%
Japan {CJK}UTF8min日本 9363 99.30% 96%
Australia australia 6522 99.13% 92%
India india 6649 99.25% 97%
@@
4.1.1 Future Work
@@
If we accepted a more probabilistic approach to video collection, we could have collected many more views for many fewer requests. However, this would have meant a less complete slice of TikTok. If from the data we have gathered, we can determine that ID patterns with more frequent video hits are not meaningfully different than ID patterns with less frequent occurrence of video posts, we could confidently roll out such a system.
@@ 4.4.2 Classification of videos featuring children
Given the difficulty of detecting children of varying age groups in contexts, the obtained F1 score of 0.86 seems particularly dependable. As stated by Schirmer et al. [56], the range of contexts in which videos feature children is vast: Videos contain vlogs of new mothers showing their newborn, parents playing with their toddlers, and families performing popular TikTok dances. Further, children might be shown only for a couple of seconds or throughout the whole video. Therefore, textual descriptions of videos can contain a lot of noise or might even miss child-related cues, making the performance of our classifier even more promising.
@@ 5 Data Availability
We have released aggregated statistics from our collected dataset in order to preserve the privacy of those in our dataset. We have released the data here URL.
@@ 6 Code Availability
The code for this work is available at URL
[…]
Supplementary information
If your article has accompanying supplementary file/s please state so here.
Authors reporting data from electrophoretic gels and blots should supply the full unprocessed scans for key as part of their Supplementary information. This may be requested by the editorial team/s if it is missing.
Please refer to Journal-level guidance for any specific requirements.

### v3

@@ 2.5 Inauthentic Activity
With any social media dataset, understanding human behaviour through it also requires understanding how much of the data is produced through inauthentic activity [28].
We know that bots exist on the platform [29], but there are no publicly available bot detection tools for TikTok. Furthermore, it is unlikely that existing bot detection methods transfer well [13], due to the difference in text distribution and platform properties. In this dataset, we do not have full chronological data timelines to allow methods that exploit timestamps to work fully [30], or network data to allow graph methods to work [31].
Given these constraints, we used simple heuristics for bot detection [32]. We find that there are 46 accounts with more than 100 videos each posted in the hour we looked at, and the top 5 highest frequency posting accounts, with between 1500 and 6000 videos in the hour we looked at, all have the default username of user followed by 9 random digits. Their videos have an average of 1.9 views each, but account for 0.28% of the total 1 hour dataset. Looking at comments, there are 831 accounts that have at least 10 comments and whose mean commenting frequency is higher than 1 every 3 seconds, with most of their comments (88%) consisting only of emojis. It’s clear from these initial statistics that there is automated activity on the platform, but these methods can only produce a rough lower bound on the extent of automated activity on the platform. We therefore hope future work can narrow our bounds on the amount of inauthentic activity on the platform and in this dataset.
@@ 3 Discussion
The data we report in this work is critical for calibrating our understanding of work on TikTok. Here we briefly illustrate the utility of the data with five exploratory case studies.
The invasion of Ukraine was touted as the ‘First TikTok War’ [33], and the way the invasion of Ukraine was covered on TikTok was investigated [34] as eyes turned towards the conflict. With this dataset, we can learn that while the average view count of posts from Ukraine (4.1k) was higher than the global average (2.2k), it is still lower than the average view count of posts from the US (6.3k), quantifying our understanding of its global attention on the day posts for this work were collected. We can also show that, in keeping with TikTok policy, no posts are tagged with a Russian origin (although of course workarounds such as VPNs for posting from Russia are possible).
In other places it’s been established that there is a tendancy in the global south towards engagement with content and culture from the global north [35]. Our data renders this in striking relief when we consider that the global south [36] accounts for 81% of video posts but only receive average view counts of 1.5k (compared to a baseline of 2.2k).
@@
Finally, for those looking at topical issues on TikTok, our data provides an entrypoint by way of topical clusters. For work on Palestinian activism on TikTok [37], we can aid in initial investigations by showing that 410 videos in our hour dataset were clustered as a Palestine protest associated topic, but those videos had a significantly lower (p < 0.02) average view count of 440 views compared to the hour dataset view average of 2.5k. For work looking at the role of war on TikTok[38], we can show that 2,995 videos in our hour dataset were clustered as a military/soldier associated topic, and those videos received an average of 5.2k views. And for work looking at the link between TikTok and public health [39], we can add context by showing that 300 videos associated with the topic ‘Medical Examination and Healthcare Setting’ had a significantly higher (p < 0.02) average view mean of 28.8k views than the average.
@@
In order to allow others to use the data we have collected here to calibrate their studies, we have released aggregate data distributions. We understand that this release won’t cover every use case, so we invite interested parties to contact us for additional specific statistics.
@@
This dataset is not perfect. Collecting the data for this dataset took 5 months, meaning that the first batch of data was collected 2 months after the day chosen, and the last batch of data was collected 7 months after the day chosen. This introduces temporal bias into our dataset, where older videos will accumulate more engagement, and are also more likely to be taken down, or deleted. However, looking at our comment engagement offset (Fig. 6), we can see that the median time of engagement for all comment authors is around 29.9 hours after the video was posted. It seems likely that the median time from video post to comment engagement is similar to the median time offset for other engagement. This means that the vast majority of engagement with posts happens within two months, meaning that at least for engagement, the temporal bias of a collection process should have minimal effect. We also know that from looking at deletion rates at different minutes in our collection, there was an 8% percentage point increase in deletion rates from the first minute of our hour to the last minute.
@@
We also cannot speak to the proportion of authentic to inauthentic content in this dataset. Detecting bots and influence operations is a pressing issue on TikTok which inhibits researchers’ abilities to study human vs. bot behavior on the platform. While we used simple heuristics to examine the least sophisticated types of bots on the platform, in the age of LLMs and advanced influence operations, more comprehensive methods must be used to be able to detect this activity, which we will leave for future work.
Despite these and other limitations, we believe the statistics we have presented here to be of substantive value to the research community. With this work we have presented the first large vertical slice of posts from TikTok, and delivered the essential statistics to help understand the platform. Our data release will allow those doing studies of TikTok to contextualize their findings, and work to deepen our understanding of this global platform.
@@ 4.1 ID Generation
We have a starting point, which is that we know that if we convert the integer to a 64 bit binary representation, we can take the first 32 bits, convert these 32 bits back to an integer representation, we get a unix timestamp, which tends to be within seconds of the recorded creation time of the video recorded in the video metadata [40] (Although for reasons we detail later, we believe this is specifically the time at which a post is created in the system, not when it is released, an important difference for scheduled posts, see Section 4.1). This indicates that within the total 64 bits of the ID, there is encoded information. Many large scale distributed systems that need to generate globally unique IDs, use systems similar to the Snowflake system [41, 42, 43, 44]. If the ID generation method is similar to snowflake, then we would expect that the rest of the bits are some combination of milliseconds data, and data relating to TikTok’s infrastructure (worker IDs, sharding IDs, sequence IDs etc.). If we can understand what composes the rest of these bits, when generating IDs, we can reduce the search space from $4.3\times 10^{9}$ combinations down to a manageable number.
@@
Canada Seedlist 118,994 99.93%
Germany Seedlist 35,256 99.94%
Nigeria Country Name Search 41,127 99.86%
Indonesia Country Name Search 19,509 99.83%
Brazil Country Name Search 14,079 99.77%
All 225,127 99.97%
@@
Due to there being datacentre specific sections of the ID, ensuring we have a good ID generation system is essential to having a globally representative collection of posts.
The Good-Turing estimator gives us confidence in what percentage of videos we have for each geographical area, but the question remains if we have good geographic coverage. After looking at the geographic distribution of the videos collected in the one minute from each hour dataset, we saw that the notable places where video counts were lower than others were East Asia, Oceania, India, and Russia. Though there are good reasons to believe that video counts in India and Russia would naturally be lower due to the ban of TikTok in those countries [24, 50], we collected extra datasets correlated with those regions (India, Australia, Japan, Russia) to check if there were previously undiscovered IDs, see details in Table 4. We found that the unique IDs we had previously collected covered at least 99% of the videos from those areas, and for India and Japan, at least 99.25%, which we feel is sufficient to ensure a dataset is representative of those areas.
@@
Russia Россия 10,288 99.35% 91%
Japan {CJK}UTF8min日本 9,363 99.30% 96%
Australia australia 6,522 99.13% 92%
India india 6,649 99.25% 97%
@@
Time Distribution Discrepancies
We noticed that when we look at post time distribution over each second of the minute using the createTime attribute in the metadata, there is a 14% increase at the $0^{th}$ second compared to the rest of the seconds, but that this distribution is flat when looking at the create time of the post as derived from the ID. As detailed in Section 2.1.1, these videos at the $0_{th}$ second are significantly different from posts in other seconds. We hypothesize that this means that the create time in the ID is the time when a scheduled post is created, and the createTime attribute is the time when the video is made available as scheduled. This would imply that some of the posts we request are posts that are not yet available, and as such, would have an error message detailing that, which we hypothesize to be the status_self_see error message.
Future Work
@@
If we accepted a more probabilistic approach to video collection, we could have collected many more views for many fewer requests. However, this would have meant a less complete slice of TikTok. If from the data we have gathered, we can determine that ID patterns with more frequent video hits are not meaningfully different than ID patterns with less frequent occurrence of video posts, we could confidently roll out such a system. Figure 12 shows the percentage of videos we can expect to get by aiming for higher hit rates.
Figure 12: Success rate of ID generation vs percentage of videos successfully fetched out of all video we find for our combined dataset.
@@ 4.4.2 Classification of videos featuring children
Given the difficulty of detecting children of varying age groups in contexts, the obtained F1 score of 0.86 seems particularly dependable. We looked at false positives and true negative predictions from the classifier, and found no general trend based on demographics. As stated by Schirmer et al. [58], the range of contexts in which videos feature children is vast: Videos contain vlogs of new mothers showing their newborn, parents playing with their toddlers, and families performing popular TikTok dances. Further, children might be shown only for a couple of seconds or throughout the whole video. Therefore, textual descriptions of videos can contain a lot of noise or might even miss child-related cues, making the performance of our classifier even more promising.
@@ 5 Data Availability
We have released aggregated statistics from our collected dataset in order to preserve the privacy of those in our dataset. We have released the data at https://doi.org/10.5281/zenodo.15330827.
@@ 6 Code Availability
The code for this work is available at https://github.com/bendavidsteel/tiktok-slice